# TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models

### v2

@@ 4 Discussion & Related Work
ARF [83] Random Forest $p({\mathbf{x}}\mid y)$
GOGGLE [47] VAE $p({\mathbf{x}}\mid y)$
@@
Limitations and Future Work. TabEBM is a general method that relies on an underlying binary classifier, and as such, its strengths and weaknesses are directly tied to this classifier. We used TabPFN because it is a well-established open-source pre-trained model for tabular data. Therefore, TabEBM inherits some of TabPFN’s limitations, particularly in scaling to a larger number of features. TabEBM can handle datasets with over 1000 samples, overcoming TabPFN’s limitation, as it processes one class at a time. In Section D.5.3, we show that TabEBM outperforms other generators on larger datasets, though the performance gains decrease as the sample size increases. Although we implement TabEBM with TabPFN in this paper, we stress that TabEBM is compatible with any classifier that can be adapted into EBMs, as described in Section 2. As foundational models for tabular data evolve [81], new models capable of handling more features and samples are expected. Integrating them into TabEBM will enhance its ability to manage high-dimensional datasets, increasing its versatility and utility. In addition, we would like to point out the imperfection of widely adopted univariate statistical fidelity metrics (i.e., Inverse KL, KS test and $\chi^{2}$ test). However, evaluating generators’ ability to capture the joining feature relationships remains an open research question [78]. We leave this for future work to explore.
@@ 5 Conclusion
We introduce TabEBM, the first tabular data augmentation method that creates class-specific EBM generators, learning the marginal distribution for each class separately. We also provide the first comprehensive analysis of tabular data augmentation across various dataset sizes. Our results demonstrate that TabEBM improves downstream performance through data augmentation on real-world datasets, outperforming other benchmark generators. The statistical evaluation confirms that TabEBM generates high-fidelity synthetic data, particularly for small datasets. We release our method as an open-source library, allowing users to generate data immediately without additional training.
@@
The authors would like to thank Francisco Vargas, Randall Balestriero, and Otilia Stretcu for their insightful discussions and valuable input early in the project. NS acknowledges the support of the U.S. Army Medical Research and Development Command of the Department of Defense; through the FY22 Breast Cancer Research Program of the Congressionally Directed Medical Research Programs, Clinical Research Extension Award GRANT13769713. Opinions, interpretations, conclusions, and recommendations are those of the authors and are not necessarily endorsed by the Department of Defense.
Appendix: TabEBM: A Tabular Data Augmentation Method with Distinct Class-Specific Energy-Based Models
[…]
@@ 3. Answer: [Yes]
  4. Justification: Please refer to Section 1, where we state our research topic and objective and then connect them to our contributions. We also provide an open-source library of the proposed method (https://github.com/andreimargeloiu/TabEBM).
@@ 7. Answer: [Yes]
  8. Justification: Please refer to Section 4, where we analyse the limitations of the proposed model and promising directions for future research.
@@ 11. Answer: [Yes]
  12. Justification: Please refer to Section 2, where we provide full theoretical derivations of our proposed method.
@@ 15. Answer: [Yes]
  16. Justification: Please refer to Appendix B, where we provide full details on reproducing the results in the paper. Furthermore, we have also provided the open-source implementation of our proposed method.
@@ 19. Answer: [Yes]
  20. Justification: Please refer to Appendix B. All datasets used in this paper are publicly available, and the implementations of benchmark generators are open-source. In addition, we have also provided the implementation of our proposed method.
@@ 23. Answer: [Yes]
  24. Justification: Please refer to Appendix B, where we provide full descriptions of the experimental setup to reproduce the results in the paper.
@@ 27. Answer: [Yes]
  28. Justification: Please refer to Section 3, where we provide standard deviations for all tables. Figure 6 and Figure 4 (Right) contain error bars. Due to the page limit, the error bars for all other figures are available in Appendix D. In Section 3.2 and Section D.6, we show statistical significance tests of the similarity between real data and synthetic data.
@@ 31. Answer: [Yes]
  32. Justification: Please refer to Section B.4, where we provide full details on the computation resources used in the paper.
@@ 39. Answer: [Yes]
  40. Justification: Please refer to Appendix A, where we include the societal impacts of our work.
@@ 47. Answer: [Yes]
  48. Justification: Please refer to Appendix B, where we provide the open-source licenses followed by the creators or original owners of assets.

### v3

@@ 4 Discussion & Related Work
ARF [83] Random Forest $p({\mathbf{x}},y)$
GOGGLE [47] GNN $p({\mathbf{x}}\mid y)$
@@
Limitations and Future Work. TabEBM is a general method that relies on an underlying binary classifier, and as such, its strengths and weaknesses are directly tied to this classifier. We used TabPFN because it is a well-established open-source pre-trained model for tabular data. Therefore, TabEBM inherits some of TabPFN’s limitations, particularly in scaling to a larger number of features. TabEBM can handle datasets with over 1000 samples, overcoming TabPFN’s limitation, as it processes one class at a time. In Section D.5.3, we show that TabEBM outperforms other generators on larger datasets, though the performance gains decrease as the sample size increases. Although we implement TabEBM with TabPFN in this paper, we stress that TabEBM is compatible with any classifier that can be adapted into EBMs, as described in Section 2. As foundational models for tabular data evolve [81], new models capable of handling more features and samples are expected. Integrating them into TabEBM will enhance its ability to manage high-dimensional datasets, increasing its versatility and utility. Finally, note that, generators that are limited in modelling multivariate distributions may still perform well on univariate fidelity metrics, which is a standard approach to evaluating such models. However, evaluating their ability to learn more complex, high-order, relationships between features remains an open research question [78], which we leave for future work.
@@ 5 Conclusion
We introduced TabEBM, the first tabular data augmentation method that creates class-specific EBM generators, learning the marginal distribution for each class separately. We also provide the first comprehensive analysis of tabular data augmentation across various dataset sizes. Our results demonstrate that TabEBM improves downstream performance through data augmentation on real-world datasets, outperforming other benchmark generators. The statistical evaluation confirms that TabEBM generates high-fidelity synthetic data, particularly for small datasets. We release our method as an open-source library, allowing users to generate data immediately without additional training.
@@
The authors would like to thank Francisco Vargas, Randall Balestriero, and Otilia Stretcu for their insightful discussions and valuable input early in the project. NS and MJ acknowledge the support of the U.S. Army Medical Research and Development Command of the Department of Defense; through the FY22 Breast Cancer Research Program of the Congressionally Directed Medical Research Programs, Clinical Research Extension Award GRANT13769713. Opinions, interpretations, conclusions, and recommendations are those of the authors and are not necessarily endorsed by the Department of Defense.
Appendix
@@ 3. Answer: [Yes]
  4. Justification: Section 1 details our research objectives and highlights our contributions.
@@ 7. Answer: [Yes]
  8. Justification: Presented in Section 4.
@@ 11. Answer: [Yes]
  12. Justification: Section 2 presents the theoretical results of our proposed method.
@@ 15. Answer: [Yes]
  16. Justification: Refer to Appendix B, where we provide full details on reproducing the results in the paper. We provide an open-source library of the proposed method.
@@ 19. Answer: [Yes]
  20. Justification: Refer to Appendix B. All datasets used in this paper are publicly available, and the implementations of benchmark generators are open-source. We also provide an open-source library https://github.com/andreimargeloiu/TabEBM .
@@ 23. Answer: [Yes]
  24. Justification: Appendix B provides full descriptions of the experimental setup.
@@ 27. Answer: [Yes]
  28. Justification: Refer to Section 3, where we provide standard deviations for all tables. Figure 6 and Figure 4 (Right) contain error bars. Due to the page limit, the error bars for all other figures are available in Appendix D. In Section 3.2 and Section D.6, we show statistical significance tests of the similarity between real data and synthetic data.
@@ 31. Answer: [Yes]
  32. Justification: Refer to Section B.4, where we provide full details on the computation resources used in the paper.
@@ 39. Answer: [Yes]
  40. Justification: Refer to Appendix A, where we include the societal impacts of our work.
@@ 47. Answer: [Yes]
  48. Justification: Refer to Appendix B, where we provide the open-source licenses followed by the creators or original owners of assets.